# Knapsack Optimization-based Schema Linking for LLM-based Text-to-SQL Generation

### v1

@@ 4.1 Relevance Estimation
Binary scoring function.
This component is tasked with providing binary relevance assessments, determining whether a schema element is relevant or not in a straightforward manner:
@@
Considering efficiency, we finetune a lightweight LLM, DeepSeek-Coder-1.3B (Guo et al., 2024), using LoRA (Hu et al., 2021), a well-established parameter-efficient technique. We employ a strategy whereby the initial generation of a simulated SQL provides a contextual foundation for the subsequent generation of schema linking. The loss function during fine-tuning is defined as follows:
@@
where $Q$ is the set of queries in the training dataset. ${\text{SQL}}^{*}_{q}$ and ${\bm{S}}^{*}$ are the ground truth SQL and linking result, respectively. This formulation utilizes a joint generation of a simulated SQL and the schema linking result, thereby ensuring that the linking generation benefits from the contextual alignment provided by the simulated SQL during both fine-tuning and inference. Noted that the simulated SQL generated during this process will not be utilized in subsequent steps. The input and output examples are provided in Table 6 in Appendix B.
Probabilistic scoring function.
In conjunction with the binary scoring model, we introduce a probabilistic scoring model to enhance prediction robustness. The primary motivation is to mitigate the risk of the binary model misclassifing relevant elements. The probabilistic scoring model assigns a soft score to each element as the linking probability to reduce the likelihood of relevant elements missing , thereby supporting the binary model:
@@
where $r^{p}_{i}\in[0,1]$ represents the probabilistic score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$.
For the probabilistic scoring model, we utilize an encoding model, RoBERTa-Large (Liu et al., 2019), to derive the semantic embeddings for the query and all schema elements. Following the methodology of (Li et al., 2023a, 2024b), we employ a cross-attention network to jointly embed the semantic representations of tables and their columns. The dot product of the query embedding and the element embedding yields the final probabilistic score. This probabilistic model is trained using the presence of $s$ in ${\bm{S}}^{*}$ as the learning objective. The loss function is defined as follows:
@@
where $\text{FL}(\cdot)$ denotes the focal loss function (Ross and Dollár, 2017), designed to emphasize learning from hard negative samples.
Relevance scoring estimation.
@@
where $r_{i}$ denotes the relevance score assigned to the $i$-th schema element $s_{i}\in{\mathcal{S}}_{q}$ given the query $q$. An upper limit of 1 is imposed on $r_{i}$ to prevent excessively strong contrasts between potentially relevant elements, as each relevant schema element is equally important for accurate SQL generation.
4.2 Redundancy Estimation
@@
Redundancy scoring estimation.
@@
$w_{i}={r_{i}}^{-1}.$ (11)
@@
It ensures elements with lower relevance scores contribute more to redundancy, thereby effectively prioritizing more relevant schema elements.
Upper Redundancy Tolerance.
@@
Redundancy tolerance defines the limit on the total redundancy score of elements in the schema linking results, as an excess of redundant elements can significantly disrupt subsequent SQL generation. Estimating it during inference is challenging. For the SQL generation of some queries, a lenient tolerance may be beneficial since it allowing the inclusion of more potentially relevant elements in the linking results to prevent missing. However, for other queries, a lenient tolerance could introduce to many redundant elements, which may significantly disrupt the final SQL generation process.
@@
To address this challenge, we hypothesize that similar queries exhibit comparable tolerance levels for redundancy and leverage queries from the training dataset, where ground truth linking results are available, to aid in predicting redundancy tolerance for given query $q$ during inference. We first retrieve the top-$K$ queries in the training dataset that closely resemble $q$ using sentence similarity measures (Gao et al., 2021). For a query $q_{k}$ in the retrieved top-$K$ queries, we calculate the total redundancy score for its ground truth linking results ${\mathcal{S}}^{*}_{q_{k}}$ and use it as its upper redundancy tolerance. The maximum of these upper redundancy tolerances from the $K$ most similar queries in the training dataset is then used as the redundancy tolerance estimate for query $q$:
$u_{q}=\max_{q_{k}\in\text{TopK}(q)}\bigg{(}\sum_{s_{j}\in{\mathcal{S}}^{*}_{q_{k}}}w_{j}\bigg{)}$ (12)
@@
By capturing the maximum sum from the most similar queries, this approach effectively defines an upper bound on redundancy tolerance that is both efficient and adaptable, ensuring that while redundant elements are excluded, the risk of missing relevant elements is minimized.
4.3 Schema Linking with Knapsack Optimization
@@
Given the relevance score $r_{i}$, the redundancy score $w_{i}$ for each $s_{i}\in{\mathcal{S}}_{q}$, and the upper redundancy tolerance $u_{q}$ for query $q$, we introduce a hierarchical schema linking strategy with knapsack optimization. This approach is designed to ensure the linking of relevant schema elements to avoid missing and minimize the inclusion of redundant ones for a precise schema linking.
Knapsack Optimization
@@
The 0-1 knapsack problem (Fréville, 2004) is a classic combinatorial optimization challenge. It involves selecting items from a finite set, each item characterized by a specific value and weight, to place into a knapsack with a limited weight capacity. The objective of 0-1 knapsack optimization is to choose items that maximize the total value while ensuring the total weight remains below the knapsack’s capacity.
@@
$\begin{split}f_{\text{Knap}}:&(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q}=\arg\max\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}r_{i}\\ &\text{s.t.}\quad\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}w_{i}\leq u_{q}.\end{split}$ (13)
This formulation aims to maximize the total relevance scores of the linked schema elements. The constraint ensures that the cumulative redundancy score remains within the upper redundancy tolerance of query $q$. To tackle this optimization problem and achieve the optimal linking result, We employ an efficient dynamic programming approach. Detailed pseudocode of this dynamic programming algorithm is provided in Algorithm A in Appendix.
Hierarchical Linking Strategy
For efficiency, our KaSLA employs a hierarchical process in schema linking: initially linking tables and subsequently linking columns within each selected table. This strategy effectively reduces the dimensionality of candidate columns, enhancing KaSLA’s capacity to handle large-scale schema linking and text-to-SQL generation tasks.
In the initial phase for a given query $q$, KaSLA aims to identify the relevant tables from the set of all available tables in schema ${\mathcal{S}}_{q}$, denoted as ${\mathcal{T}}_{q}$:
$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}(q,{\mathcal{T}}_{q}).$ (14)
After determining tables, for each $t\in\widehat{{\bm{T}}}_{q}$, let ${\mathcal{C}}^{t}_{q}$ represent the columns in table $t$. KaSLA then optimizes column linking within each selected table:
$\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}(q,{\mathcal{C}}^{t}_{q}).$ (15)
The final schema linking result $\widehat{{\bm{S}}}_{q}$ is the union of all selected tables and their corresponding selected columns, represented as:
@@
We provide a complexity analysis is detailed in Appendix D, offering insights into KaSLA’s implementation and computational efficiency.
5 Experiments
@@
We conducted comprehensive experiments to evaluate KaSLA and address the following research questions: RQ1: Does KaSLA enhance existing text-to-SQL baselines by improving their schema linking capabilities? RQ2: Does KaSLA demonstrate superior schema linking performance by handling element missing and redundancy? RQ3: Does each component in KaSLA contribute to the overall performance, and is KaSLA an efficient model for real-world text-to-SQL applications?
5.1 Experiment Setup
We conducted experiments on two well-know large-scale text-to-SQL datasets, BIRD (Li et al., 2023b) and Spider (Yu et al., 2018), incorporating multiple robust baselines based on LLMs with in-context learning and fine-tuning. For evaluating text-to-SQL, we utilized Execution Accuracy (EX) (Yu et al., 2018; Li et al., 2023b) and Valid Efficiency Score (VES) (Li et al., 2023b). We included multiple robust baselines from two categories:(i) LLMs + in-context learning baselines: C3-SQL (Dong et al., 2023), DIN-SQL (Pourreza and Rafiei, 2023), MAC-SQL (Wang et al., 2024), DAIL-SQL (Gao et al., 2024), SuperSQL (Li et al., 2024a), Dubo-SQL (Thorpe et al., 2024), TA-SQL (Qu et al., 2024), E-SQL (Caferoğlu and Ulusoy, 2024), CHESS (Talaei et al., 2024); and (ii) LLMs + fine-tuning baselines: DTS-SQL (Pourreza and Rafiei, 2024), CodeS (Li et al., 2024b).
5.2 Main Results of SQL Generation (RQ1)
To address RQ1, we conducted experiments to evaluate the SQL generation capabilities of state-of-the-art (SOTA) baselines enhanced by our KaSLA, with results presented in Table 1. KaSLA replaces the original schema linking component in models that already have schema linking capabilities or implements a full schema prompt substitution for those that do not.
The results demonstrate that incorporating KaSLA consistently improves both Execution Accuracy (EX) and Valid Efficiency Score (VES) across baselines with diverse base models, including GPT-4, CodeS-15B, StarCoder2-15B, and Deepseek-v3. Notably, in the BIRD-dev dataset, KaSLA enhances the EX of CHESS with Deepseek-v3 from 61.34% to 63.30%. For local LLMs, both "CodeS + KaSLA" and "TA-SL + KaSLA" with CodeS-15B and StarCoder-15B achieve impressive increases in Execution Accuracy compared to their original versions. These results underscore KaSLA’s effectiveness and flexibility as a plug-in, showing that KaSLA plays an important role in improving complex text-to-SQL tasks. This adaptability not only confirms its robustness but also broadens its applicability across different models, achieving superior performance outcomes and making it a versatile addition to existing systems.

### v2

@@ IV-B1 Relevance Estimation
Binary Scoring Function
This component is tasked with providing binary relevance assessments, determining whether a schema element is relevant or not in a straightforward manner with hard 0/1 relevance decisions:
@@
Considering efficiency, we finetune a lightweight LLM, DeepSeek-Coder-1.3B [14], with a well-established parameter-efficient technique, LoRA [15]. We employ a strategy whereby the initial generation of a simulated SQL provides a contextual foundation for the subsequent generation of schema linking. The loss function during fine-tuning is defined as follows:
@@
where $Q$ is the set of queries in the training dataset. ${\text{SQL}}^{*}_{q}$ and ${\bm{S}}^{*}$ are the ground truth SQL and linking result, respectively. This formulation utilizes a joint generation of a simulated SQL and the schema linking result, thereby ensuring that the linking generation benefits from the contextual alignment provided by the simulated SQL during both fine-tuning and inference. Noted that the simulated SQL generated during this process will not be utilized in subsequent steps.
Probabilistic Scoring Function
In conjunction with the binary scoring model, we introduce a probabilistic scoring model to enhance prediction robustness. The primary motivation is to mitigate the risk of the binary model misclassifing relevant elements. The probabilistic scoring model assigns a soft score to each element as the linking probability to reduce the likelihood of the missing of potential relevant elements, thereby supporting the binary model. The equation is defined as follows:
@@
where $r^{p}_{i}\in[0,1]$ represents the probabilistic score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$. A value closer to 1 indicates that the element is considered more relevant to query q.
For the probabilistic scoring model, we utilize an encoding model, RoBERTa-Large [16], to derive the semantic embeddings for the query and all schema elements. Following the methodology of  [7, 8], we employ a cross-attention network to jointly embed the semantic representations of tables and their columns. The dot product of the query embedding and the element embedding yields the final probabilistic score. This probabilistic model is trained using the presence of $s$ in ${\bm{S}}^{*}$ as the learning objective. The loss function is defined as follows:
@@
where $\text{FL}(\cdot)$ denotes the focal loss function  [17], designed to emphasize learning from hard negative samples.
Relevance Scoring Estimation
@@
where $r_{i}$ denotes the relevance score assigned to the $i$-th schema element $s_{i}\in{\mathcal{S}}_{q}$ given the query $q$. An upper limit of 1 is imposed on $r_{i}$ to prevent excessively strong contrasts between potentially relevant elements, as each relevant schema element is equally important for accurate SQL generation. We use ${{\mathcal{R}}}_{{\mathcal{S}}_{q}}=\{r_{1},\ldots,r_{|{\mathcal{S}}_{q}|}\}$ to denote the relevance score set of ${\mathcal{S}}_{q}$.
IV-B2 Redundancy Estimation
@@
Redundancy Scoring Estimation
@@
$\begin{split}f_{\text{redundancy}}:&(q,{{\mathcal{R}}}_{{\mathcal{S}}_{q}})\mapsto\{w_{1},\ldots,w_{|{\mathcal{S}}_{q}|}\}\\ &=\left\{{r_{i}}^{-1}\mid r_{i}\in{{\mathcal{R}}}_{{\mathcal{S}}_{q}}\right\}.\end{split}$ (13)
@@
It ensures elements with lower relevance scores contribute more to redundancy, thereby effectively prioritizing more relevant schema elements. We use ${{\mathcal{W}}}_{{\mathcal{S}}_{q}}=\{w_{1},\ldots,w_{|{\mathcal{S}}_{q}|}\}$ to denote the redundancy score set of ${\mathcal{S}}_{q}$.
Redundancy Tolerance
@@
Upper tolerance of redundancy defines the limit on the total redundancy score of elements in the schema linking results, as an excess of redundant elements can significantly disrupt subsequent SQL generation. Estimating a suitable upper redundancy tolerance during inference is challenging. For the SQL generation of some queries, a lenient tolerance may be beneficial since it allowing the inclusion of more potentially relevant elements in the linking results to prevent missing. However, for other queries, a lenient tolerance could introduce too many redundant elements, which might significantly disrupt the final SQL generation process.
@@
To address this challenge, we hypothesize that similar queries have comparable tolerance levels for redundancy. We sample queries from the training dataset where ground truth linking results are available, and use these paired samples to construct a query-linking pool. During inference, when predicting the redundancy tolerance for a given query $q$, we first retrieve the top-$K$ most similar queries from the query-linking pool using sentence similarity measures [18]. For a query $q_{k}$ in the retrieved top-$K$ queries, the redundancy score set for its ground truth linking results ${\mathcal{S}}^{*}_{q_{k}}$ can be denoted as ${{\mathcal{W}}}_{{\mathcal{S}}^{*}_{q_{k}}}$. We sum the total ${{\mathcal{W}}}_{{\mathcal{S}}^{*}_{q_{k}}}$ and use it as the upper redundancy tolerance of $q_{k}$. The maximum of these upper redundancy tolerances of the $K$ most similar queries from the query-linking pool is then used as the redundancy tolerance estimate for query $q$:
$f_{\text{tolerance}}:(q)\mapsto u_{q}=\max_{q_{k}\in\text{TopK}(q)}\sum_{w_{j}\in{{\mathcal{W}}}_{{\mathcal{S}}^{*}_{q_{k}}}}w_{j}.$ (14)
@@
By capturing the maximum redundancy tolerances of the most similar queries, this approach effectively defines an upper bound on redundancy tolerance that is both efficient and adaptable, ensuring that while redundant elements are excluded, the risk of missing relevant elements is minimized.
IV-C Schema Linking with Knapsack Optimization
@@
Given (i) the relevance score $r_{i}$ and redundancy score $w_{i}$ for each candidate element $s_{i}\in{\mathcal{S}}_{q}$, and (ii) the upper redundancy tolerance $u_{q}$ defined at the query level, we propose a hierarchical schema linking strategy incorporating knapsack optimization. This approach is designed to ensure the linking of relevant schema elements to avoid missing and minimize the inclusion of redundant ones for a precise schema linking.
IV-C1 Knapsack Optimization
@@
The 0-1 knapsack problem [13] is a classic combinatorial optimization problem. It involves selecting items from a finite set. Given a set of items, each with a value and a weight, the goal is to select a subset of items that maximizes the total value while ensuring the total weight does not exceed a limited knapsack capacity. In the “0-1” variant, each item is either fully included or excluded (i.e., no fractional selections are permitted).
@@
$\begin{split}f_{\text{Knap}}:&(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q}\\ &=\arg\max\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}r_{i},\quad\text{s.t.}\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}w_{i}\leq u_{q}.\end{split}$ (15)
This formulation aims to maximize the total relevance scores of the linked schema elements with a total redundancy constraint. The constraint ensures that the cumulative redundancy score remains within the upper redundancy tolerance of query $q$. To tackle this optimization problem and achieve optimal linking results, we employ an efficient dynamic programming approach. The details of this algorithm, as implemented in KaSLA, are presented in Algorithm 1.
Input: relevance score set ${\mathcal{R}}$, redundancy score set ${\mathcal{W}}$, the upper redundancy tolerance $u$
Output: schema linking set $\hat{S}$
$n\leftarrow|{\mathcal{R}}|$, $\hat{S}\leftarrow\{\}$
[…]
$keep\leftarrow\text{array of }(n+1)\times(u+1),\text{ initialized to False}$
for $i\leftarrow 1$ to $n$ do
       for $w\leftarrow 0$ to $u$ do
             if ${\mathcal{W}}[i-1]\leq w$ then
                         $keep[i][w]\leftarrow\text{True}$
                  else
                         $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
            else
                   $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
$k\leftarrow u$  for $i\leftarrow n$ downto $1$ do
       if $keep[i][k]$ then
             $\hat{S}.\text{add}(i-1)$
             $k\leftarrow k-{\mathcal{W}}[i-1]$
return $\hat{S}$
Algorithm 1 Dynamic Programming-based 0-1 Knapsack Optimization Algorithm
IV-C2 Hierarchical Linking Strategy
To ensure computational efficiency, KaSLA adopts a hierarchical schema linking strategy consisting of two stages: (1) table linking followed by (2) column linking within selected tables. This hierarchical approach significantly reduces the search space of candidate columns, enabling KaSLA to efficiently handle large-scale databases while maintaining robust performance.
In the initial phase for a given query $q$ and the corresponding ${\mathcal{S}}_{q}$, KaSLA aims to identify the relevant tables from the set of all available tables in schema ${\mathcal{S}}_{q}$, denoted as ${\mathcal{T}}_{q}$:
$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{T}}_{q}},{{\mathcal{W}}}_{{\mathcal{T}}_{q}},u^{{\mathcal{T}}}_{q}),$ (16)
where ${{\mathcal{R}}}_{{\mathcal{T}}_{q}}$ and ${{\mathcal{W}}}_{{\mathcal{T}}_{q}}$ are the sets of estimated relevance and redundancy scores for tables in schema ${\mathcal{S}}_{q}$, and $u^{{\mathcal{T}}}_{q}$ is the upper redundancy tolerance for table linking. Different from the general schema elements ${\mathcal{S}}_{q}$ used in the equations in Section IV-B, the notation ${\mathcal{T}}_{q}$ and ${\mathcal{T}}$ specifically refer to tables in schema linking. After determining the selected tables, for each $t\in\widehat{{\bm{T}}}_{q}$, let ${\mathcal{C}}^{t}_{q}$ represent the columns in table $t$. KaSLA then optimizes column linking within each selected table:
$\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}},{{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}},u^{{\mathcal{C}}}_{q}),$ (17)
where ${{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}}$ and ${{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}}$ are the sets of estimated relevance and redundancy scores for columns in table $t$, and $u^{{\mathcal{C}}}_{q}$ is the upper redundancy tolerance for column linking. The final schema linking result $\widehat{{\bm{S}}}_{q}$ is the union of all selected tables and their corresponding selected columns, represented as:
@@ IV-C2 Hierarchical Linking Strategy
We present the complete procedure of the hierarchical linking strategy in Algorithm 2.
Input: user query $q$, full schema ${\mathcal{S}}_{q}$
Output: schema linking results $\widehat{{\bm{S}}}_{q}$, including a table linking set $\widehat{{\bm{T}}}_{q}$ and the corresponding column linking set $\widehat{{\bm{C}}}^{t}_{q}$ for each table $t\in\widehat{{\bm{T}}}_{q}$
$({{\mathcal{R}}}_{{\mathcal{T}}_{q}},\left\{{{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}}\right\}_{t\in{\mathcal{T}}_{q}})=f_{\text{relevance}}(q,{\mathcal{S}}_{q})$,
${{\mathcal{W}}}_{{\mathcal{T}}_{q}}=f_{\text{redundancy}}(q,{{\mathcal{R}}}_{{\mathcal{T}}_{q}}),(u^{{\mathcal{T}}}_{q},u^{{\mathcal{C}}}_{q})=f_{\text{tolerance}}(q)$
$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{T}}_{q}},{{\mathcal{W}}}_{{\mathcal{T}}_{q}},u^{{\mathcal{T}}}_{q})$ // Table linking with Alg. 1
for $t\in\widehat{{\bm{T}}}_{q}$ do
       ${{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}}=f_{\text{redundancy}}(q,{{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}})$
       $\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}({{\mathcal{R}}}_{{\mathcal{C}}^{t}_{q}},{{\mathcal{W}}}_{{\mathcal{C}}^{t}_{q}},u^{{\mathcal{C}}}_{q})$ // Column linking with Alg. 1
return $\widehat{{\bm{S}}}_{q}=\widehat{{\bm{T}}}_{q}\cup\bigcup_{t\in\widehat{{\bm{T}}}_{q}}\widehat{{\bm{C}}}^{t}_{q}$
Algorithm 2 Hierarchical Linking Strategy
V Experiments
@@
We conducted comprehensive experiments to evaluate KaSLA and address the following research questions: RQ1: Does KaSLA enhance existing text-to-SQL baselines by improving their schema linking capabilities? RQ2: Does KaSLA demonstrate superior schema linking performance by handling element missing and redundancy? RQ3: Does each component in KaSLA contribute to the overall performance, and what is the effect of parameter scaling? RQ4: Is KaSLA an efficient model for real-world text-to-SQL applications, especially in terms of its ability to transfer to unseen domains?
V-A Experiment Setup
V-A1 Datasets and Evaluation Metrics
We conducted experiments on two well-known large-scale text-to-SQL datasets, BIRD [19] and Spider [20]. Both datasets feature human-annotated queries and SQLs, complex database elements, and challenging cross-domain scenarios meticulously. For text-to-SQL evaluation, we measured performance using two well-established execution-based metrics: Execution Accuracy (EX) [20, 19] and Valid Efficiency Score (VES) [19].
V-A2 Baselines
We incorporating multiple robust baselines based on LLMs with in-context learning and fine-tuning. We included baselines from two categories:(i) LLMs + in-context learning baselines: C3-SQL [21], DIN-SQL [22], MAC-SQL [23], DAIL-SQL [4], SuperSQL [24], Dubo-SQL [25], TA-SQL [5], E-SQL [26], CHESS [11], RSL-SQL [12]; and (ii) LLMs + fine-tuning baselines: DTS-SQL [9], TA-SQL [5], CodeS [8]. We also introduced Pure-SL, a baseline that directly fine-tunes LLMs for schema linking without any specially designed mechanisms. For baseline methods that include a schema linking module which can be cleanly separated from their main framework, we distinguish these variants by adding an ’-SL’ suffix to their original names. For example, a baseline named “CodeS” would become “CodeS-SL” when referring specifically to its schema linking component.
V-A3 Implementation Details
We implemented the proposed KaSLA and baselines under the following settings: (i) Close-source LLM-based experiments: We conducted experiments for most in-context learning baselines using GPT-4-turbo. For E-SQL, CHESS, and RSL-SQL, which originally used GPT-4o, we switched to deepseek-V3 because the API cost of GPT-4o was too high for these methods. Since previous work has shown that deepseek-V3 achieves similar performance to GPT-4o at a much lower cost [27]. (ii) Open-source LLM-based experiments: We use CodeS-15B [8] and StarCoder2-15B [28] as the base LLMs for the SQL generation model, and reproduce all baselines using LoRA [15] for parameter-efficient fine-tuning. The learning rate is set to $41e-$ 4 with cosine decay, the batch size is 16, and the training runs for 3 epochs. (iii) KaSLA: We set the top-$K$ value for redundancy tolerance estimation to 30. For the scoring models in KaSLA, we fine-tune a RoBERTa-Large [16] as the binary scoring model following [8], and a Deepseek-coder-1.3B [14] as the probabilistic scoring model using the same fine-tuning settings with all fine-tuning baselines as described in (ii).
TABLE I: The SQL generation performance of enhanced text-to-SQL models using KaSLA, which features 1.6 B parameters combining DeepSeek-coder-1.3B and RoBERTa-Large, is evaluated in terms of Execution Accuracy (EX) (%) and Valid Efficiency Score (VES) (%) on the BIRD-dev and Spider-dev datasets. We use KaSLA to replace the original schema linking component or the full schema in the baseline models, denoted as “+ KASLA”, and the relative improvements are highlighted with a gray background.